OpenStudy (anonymous):

solve the differential equation dy/dx + 2y/x =x2

sam (.sam.):

Do you know what's the integrating factor here?

sam (.sam.):

This is a first-order linear ODE

sam (.sam.):

So, \[\frac{dy}{dx}+P(x)y=Q(x)\] Integrating factor \[\huge \mu(x)=e^{\int\limits P(x)dx}\]

OpenStudy (anonymous):

you doing better

sam (.sam.):

So what's P(x)?

OpenStudy (anonymous):

2/x

sam (.sam.):

Yes so, \[\Large \mu(x)=e^{\int\limits \frac{2}{x}dx}=e^{2lnx}=x^2\] Now multiply \(\mu(x)\) across \[\frac{dy}{dx} + \frac{2y}{x} =x^2 \\ \\ x^2\frac{dy}{dx} + 2xy=x^4\] Then use reverse product rule \[\frac{d}{dx}(x^2y)=x^4\] Integrate both sides and you're done

OpenStudy (anonymous):

please show me last resultant

sam (.sam.):

Just integrate \[\int\limits \frac{d}{dx}(x^2y)dx=\int\limits x^4 dx \\ \\ x^2y=\frac{x^5}{5}+c_1\] \[y=\frac{x^3}{5}+\frac{c_1}{x^2}\]

OpenStudy (anonymous):

thank you very much

sam (.sam.):

yw